OpenStudy (anonymous):

The sum of two consecutive odd integers is 244. What is the smaller integer?

OpenStudy (zzr0ck3r):

sum

OpenStudy (zzr0ck3r):

n+(n+2) = 244

OpenStudy (zzr0ck3r):

solve for n

OpenStudy (ankit042):

sry i didn't see sum :(

OpenStudy (zzr0ck3r):

take some odd number n the next odd number is n+2 (think of 3 and 5) we need it so that when we add them we get 244 so n+(n+2) = 244

OpenStudy (ankit042):

but n can be even as well (2 and 4). To choose two odd numbers I think we should always use (2n+1) even+odd = odd

OpenStudy (anonymous):

it's 121

OpenStudy (zzr0ck3r):

good call scratch the last one let our first odd number be 2k+1 for some k let our second number be (2k+1) + 2 = 2k+3 (this is the next odd) then (2k+1)+(2k+3) = 244 4k+4 = 244 4k=240 k=60 so our first odd is 2(60) + 1 = 121